OpenStudy (anonymous):

find the slope of the line passing through (1,5) and 4,-2)

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (-2-5)/(4-1) m = (-7)/3 m = -7/3 So the slope of the line that passes through the two points is -7/3

OpenStudy (anonymous):

can you tel me how to get there and etc

OpenStudy (anonymous):

nvm i got it u just subtract

jimthompson5910 (jim_thompson5910):

The first point is (x1,y1), which is really (1,5) So (x1,y1) = (1,5) This means x1 = 1 and y1 = 5

jimthompson5910 (jim_thompson5910):

The same thing applies to the second point

OpenStudy (anonymous):

thank you.. so you just subtract basically

jimthompson5910 (jim_thompson5910):

yes, subtract the corresponding coordinates (in the same order) and reduce

OpenStudy (anonymous):

can you help me do this.. change 2x+3y = 6 into slope intercept form

jimthompson5910 (jim_thompson5910):

2x+3y = 6 3y = 6 - 2x y = (6-2x)/3 y = 6/3 - 2x/3 y = 2 - 2/3x y = -2/3x + 2

OpenStudy (anonymous):

can you work me through it on how to do it step by step

OpenStudy (anonymous):

write the equation of the line with y-intercept of 3 and slope 6

jimthompson5910 (jim_thompson5910):

The general equation is y = mx+b where m is the slope and b is the y-intercept We're given a slope of 3. So m = 3. We're also given a y-intercept of 6. So b = 6 This means that the answer is y = 3x+6